OpenStudy (anonymous):

x^2+6x=-9 solve by factoring

OpenStudy (kamille):

you need to put -9 to left side. You get: x^2+6x+9=0 Do you know formula: \[(a+b)^{2}=a ^{2}+2ab+b\]

OpenStudy (anonymous):

Yes but it still kind of confuses me :/

OpenStudy (kamille):

Do you agree, that x^2+6x+9=0 can be written as (according to the formula): \[(x+3)^{2}=0\]

OpenStudy (anonymous):

Yes

OpenStudy (kamille):

so, we have just factorised. Now we need to solve it \[(x+3)^{2}=0\] when \[x+3=0\] x=-3

OpenStudy (anonymous):

Would it be +/- 3 ?

OpenStudy (kamille):

No, it wouldnt. try to put "3" in the eq and you will get \[(3+3)^{2}=0\] \[6^{2}=0\] \[36\neq0\] And when you put "-3" in the eq, you get: \[(-3+3)^{2}=0\] \[0^{2}=0\] \[0=0\] Do you understand it?

OpenStudy (anonymous):

Yes thank you so much

OpenStudy (kamille):

You are welcome!